OpenStudy (anonymous):

Find fog and gof and their domains for the pair of functions F(x)=2x^2+3 and g(x)=4x-5 (I have fog, I cannot figure out gof)

OpenStudy (fibonaccichick666):

so first, what is g composed f?

OpenStudy (fibonaccichick666):

the function I mean,

OpenStudy (anonymous):

g should be \[2x^2+3\]

OpenStudy (fibonaccichick666):

that's what you have for f?

OpenStudy (anonymous):

g(x)=4x-5

OpenStudy (fibonaccichick666):

ok, so I am asking, when you plug f into g, what is the resulting function?

OpenStudy (fibonaccichick666):

ie. \(g(f(x) ~or~ g ~o~ f\)

OpenStudy (kc_kennylau):

Just making things tidier for everyone: \[\Large f(x)=2x^2+3\]\[\Large g(x)=4x-5\]\[\Large(f\circ g)(x) = 2(4x-5)^2+3\]\[\Large(g\circ f)(x) = 4(2x^2+3)-5\] (Please correct me if I'm wrong)

OpenStudy (anonymous):

\[32x^2-80x+53\]

OpenStudy (anonymous):

that was fog

OpenStudy (anonymous):

i dont know what gof is...I need help with that I only know fog, which I just gave an answer of

OpenStudy (fibonaccichick666):

how did you get fog, what did you do? (gof is the same process)

OpenStudy (fibonaccichick666):

or ok, let's say let x=2, what would you do?

OpenStudy (fibonaccichick666):

you'd put a 2 everywhere you see an x right?

OpenStudy (fibonaccichick666):

so in the equation g(x) replace every x with f.

OpenStudy (fibonaccichick666):

then replace every f with what f equals, it's just substitution

OpenStudy (anonymous):

I did \[fog=f(4x-5)\] \[fog=2(4x-5)^2+3\] which equals \[32x^2-80x+53\]

OpenStudy (anonymous):

that answer is for fog and now I don't know how to do gof

OpenStudy (usukidoll):

for g o f you place your f(x) function into the g(x) function

OpenStudy (fibonaccichick666):

same exact way as above

OpenStudy (fibonaccichick666):

you did it correctly

OpenStudy (fibonaccichick666):

its just g(f) instead of f(g)

OpenStudy (fibonaccichick666):

do you see what I mean?

OpenStudy (usukidoll):

some people use different terminology symbols... I don't use that but I do use g o f a lot... read this from right to left... your f(x) function should be placed inside the g(x) function

OpenStudy (anonymous):

I did and it keeps telling me I'm wrong. This is how I would do it... \[4(32x^2-80x+53)-5\] which would give me \[128x^2-320x+207\]

OpenStudy (fibonaccichick666):

ah, you put the wrong function in

OpenStudy (anonymous):

how do I put it in :(

OpenStudy (fibonaccichick666):

you put your answer for part a in instead of doing the new problem

OpenStudy (fibonaccichick666):

you plugged f composed with g in instead of just f

OpenStudy (usukidoll):

do you see the x in g(x) = 4x- 5 replace that x in g(x) with that entire f(x) function

OpenStudy (fibonaccichick666):

so what is f? Plug that in instead of your answer for part a, then I will double check it.

OpenStudy (fibonaccichick666):

you did set it up correctly, just misread your paper.

OpenStudy (anonymous):

ok so \[g(x)=4(2x^2+3)-5\]

OpenStudy (fibonaccichick666):

much better :)

OpenStudy (usukidoll):

yes so all you have to do is distribute that 4 for 2x^2+3

OpenStudy (fibonaccichick666):

go ahead, try and simplify

OpenStudy (anonymous):

final answer should be \[8x^2+7\] right?

OpenStudy (usukidoll):

we are multiplying

OpenStudy (fibonaccichick666):

well, how do you distribute the 4?

OpenStudy (usukidoll):

so what is 4 times 2x^2 and what is 4 times 3

OpenStudy (fibonaccichick666):

so just remember, if you have \[a(c+b)=ac+ab\] You have to send the a to b and c

OpenStudy (anonymous):

8x^2+12

OpenStudy (anonymous):

\[8x^2+12x-5\]

OpenStudy (fibonaccichick666):

ok now that is what happens when you distribute correctly, now, you still had that - five hanging out

OpenStudy (usukidoll):

we don't need the x for the 12

OpenStudy (usukidoll):

\[g(x)=8x^2+12-5\] what is 12 - 5 ?

OpenStudy (anonymous):

so its \[8x^2+7\]

OpenStudy (fibonaccichick666):

yea

OpenStudy (fibonaccichick666):

You were right to begin with, I apologize

OpenStudy (fibonaccichick666):

I got lost in usuki's comments

OpenStudy (anonymous):

Oh thanks...yeah I was confused that you guys said it was wrong haha

OpenStudy (fibonaccichick666):

yea, I'm sorry about that. I read the we are multip[lying and assumed you added, I didn't notice they come out the same. But anyways, does this make sense now?

OpenStudy (anonymous):

Yeah, is that the final answer?

OpenStudy (usukidoll):

yeah

OpenStudy (fibonaccichick666):

well, yes and no

OpenStudy (usukidoll):

how is this no?

OpenStudy (fibonaccichick666):

so that is the final answer for what is g composed f, but we are not just loooking for that

OpenStudy (fibonaccichick666):

now, we need to find their domain, for both bantaray

OpenStudy (anonymous):

we are looking for domains

OpenStudy (fibonaccichick666):

now, what does domain mean bant?

OpenStudy (usukidoll):

oh geez total miss. . .

OpenStudy (fibonaccichick666):

I got this usuki, you can go

OpenStudy (anonymous):

the domain for fog is \[(-\infty,\infty)\]

OpenStudy (fibonaccichick666):

h/o I need to find what that was again haha

OpenStudy (fibonaccichick666):

sounds right though

OpenStudy (fibonaccichick666):

yep, that's right

OpenStudy (fibonaccichick666):

now for g?

OpenStudy (fibonaccichick666):

side note, you can also say All real numbers

OpenStudy (anonymous):

Ok, I got the domains now...I figured them out really fast :) I think we are done

OpenStudy (fibonaccichick666):

yep, what did you get for g(f)?

OpenStudy (anonymous):

I have one more question.....Would you mind staying to help me?

OpenStudy (fibonaccichick666):

can you pop open another question and tag me?

OpenStudy (fibonaccichick666):

just close this one to make a new one. It's a bit of a long thread

OpenStudy (anonymous):

yeah sure!